How many phosphodiester bonds are present in a 1000bp plasmid?
Answer #1

plasmid double stranded DNA molecule it is 1000bp long that means it has 1000 + 1000=2000 nucleotide

phosphodiestercind can be calculate by formula

total number of phosphodiester bonds = Total number of nucleotides - 2

= 2000-2= 1998

1998 phosphodiester bonds are present in a 1000bp plasmid